OpenStudy (anonymous):

Bailey's rectangular dog pen for his Irish setter must have an area of 300 square feet. Also, the length must be 30 feet longer than the width. Find the dimensions of the pen. Working on solving using quadratic formula.

OpenStudy (emilylauren):

still need help?

OpenStudy (emilylauren):

Area =length x width

OpenStudy (emilylauren):

the problem tells us that length= width + 30

OpenStudy (emilylauren):

300= Area (w+30)w=300 so 300= w^2 +30w

OpenStudy (emilylauren):

w^2+30w-300=0 this is your quadratic equation

OpenStudy (emilylauren):

w=7.9128 ABOUT this isn't exact if it needs to be exact then \[-15+5\sqrt{21}\]

OpenStudy (emilylauren):

that is your width

OpenStudy (emilylauren):

\[15+5\sqrt{21}\] is your length

OpenStudy (emilylauren):

I wouldn't mind a medal...